OpenStudy (anonymous):

PLEASE HELP: YOU FIND 93 DINOSAUR BONES AND A COMPLETE DINOSAUR HAS 465 BONES WHAT PERCENT OF BONES HAVE YOU FOUND?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

93 out of 465 93/465=31/155

OpenStudy (anonymous):

31/155=1/5=20/100 so 20%

OpenStudy (anonymous):

thank you!!!! \

OpenStudy (anonymous):

NOW THE NEXT ONE IS: THEN YOU FIND 186 BONES OUT OF 465 BONES WHAT PERCENT IS STILL MISSING?

OpenStudy (anonymous):

Well for this problem you can set up a proportion. I use %/100 = is/of. So we know there are 93 bones of 465. this is 93/465. so we plug in our equation %/100=93/465..... %(465)=100(93). 100(93)= 9300.... soo %(465)=9300... to get % by itself, we take 9,300 divided by 465 which equals 20. %=20

OpenStudy (anonymous):

same thing. ^ 40%